生産加工特論 2 単位 (選択)

Advanced Production Technology  石田徹 · 教授/ 大学院ソシオテクノサイエンス研究部

多田吉宏 · 准教授/ 知的力学システム工学専攻機械創造システム工学コース生産システム講座

【授業目的】粉末冶金技術を中心に，生産加工論を学ぶ． 【授業概要】各種多機能な新素材に対する成形加工技術，粉末� 粉体の加工とそ れを応用した材料創製技術など新しい生産加工論を講義するとともに，圧縮 性材料の力学的取り扱いを解説する． 【授業形式】講義 【キーワード】粉末冶金，多孔質金属，塑性力学，構成式 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【履修要件】塑性加工について基礎的な素養があること． 【履修上の注意】 [注意] 【到達目標】 1. 粉末成形や高静水圧加工法などの粉末冶金法技術について理解する． 2. 圧縮性材料の塑性力学の基礎を理解する． 【授業計画】 1. 粉末冶金法の概論 2. 粉体の成形法 1 3. 粉体の成形法 2 4. 粉体の成形法 3 5. 粉体の成形性の評価法 6. 各種粉体の成形特性 7. 焼結体の加工性 8. 演習 9. 圧縮性材料の降伏条件 10. 圧縮性材料の塑性力学 11. 圧縮性材料の加工の解析 12. 粉粒体の成形条件 13. 圧縮性材料に対する変分原理 14. 圧縮性材料に対する上界定理 15. まとめ 16. 定期試験 【成績評価基準】レポート課題 (70 点) と演習・試験 (30 点) を総合して評価する．【教科書】プリント資料で講義する．【参考書】粉体の成形と加工，コロナ社， ISBN4-339-04367-2

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216727 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 石田(M321， 088-656-7379， ishidat@me.tokushima-u.ac.jp) ⇒ 多田(M319， 088-656-7381， tada@me.tokushima-u.ac.jp)

【備考】授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習を行うことが，授業の理解と単位取得のために必要である

Advanced Production Technology 2 units (selection)

Toru Ishida · Professor / Institute of Technology and Science

, Yoshihiro Tada · Associate Professor / Production Systems Engineering, Mechanical Engineering, Intelligent Structures and Mechanics Systems Engineering

Target⟩ This class introduce powder processing and plasticity theory that can be

used for novel material fabrication and advanced manufacturing. Outline⟩ Forming technologies for novel materials are introduced, and plasticity

theories for compressive metals (porous metals) are also discussed. Style⟩ Lecture �������⟩ powder metallurgy, porous metal, plasticity theory, constitutive

equation Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ Students are required to have a good understanding on fundamentals

of metal forming technologies. Notice⟩ [注意]

Goal⟩

1. To understand advanced forming technologies including powder processings and isostatic processings.

2. To understand fundamentals of compressive plasticity theory.

Schedule⟩

1. Introduction to powder metallurgy

2. powder forming processings 1

3. powder forming processings 2

4. powder forming processings 3

5. testing methods for consolidation characteristics of powder

6. consolidation characteristics of powder

7. formability of sintered preform

8. exercise

9. yield criterion of porous metal

10. plasticity theory for porous metal

11. analysis of plastic deformation of porous metal

12. consolidation criterion of powder

13. variational principle for porous metal

14. upper bound theorem for porous metal

15. conclusion

16. examination Evaluation Criteria⟩ Assignments count 70%, exercises and examinations count

30%. Textbook⟩ Printed synopsises are used. Reference⟩ Advances in Powder Forming Processes and Related Technologies,

Corona Publ.Co.,ISBN4-339-04367-2

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216727

Student⟩ Able to be taken by only specified class(es)

Contact⟩ ⇒ 石田(M321， 088-656-7379， ishidat@me.tokushima-u.ac.jp) ⇒ Tada(M319 088-656-7381 tada@me.tokushima-u.ac.jp)

Note⟩ Students are required preparation (2 hours) and review (2 hours) for every two hours class work.